OpenStudy (anonymous):

Evaluate the following if a = 2, b = -3, and c = -4 (6a-b) / (c^2)

OpenStudy (anonymous):

hold up i am figuring it out

OpenStudy (anonymous):

((6x2) - (-3) / (-4)^2

OpenStudy (anonymous):

fist say what does (-4)^2 equal to?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

correct :) 6x2 = ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

i put the answer in my algebra calculator and it came out -15/16

OpenStudy (anonymous):

@Jamie_Elizabeth , in openstudy , DON'T GIVE ANSWER!

OpenStudy (anonymous):

yes @pretty_not_smart , it's 12 and say this too: -(-3) = (-1) x (-3) = ?

OpenStudy (kewlgeek555):

Okay, so the equation started out like this:\[(6a-b)\div(c^2)\] But now the equation is turned into this: \[(6\times2--3)\div(-4^2)\] Lets see. You have to think PEMDAS. Now...here it is step-by-step. Since there are two paranthesis, it might SEEM confusing, but it really isn't. You just have to remember to work left to right.\[6\times2=12\] That makes our equation:\[(12--3)\div(-4^2)\] Now...\[12--3=15\] That makes our equation look like this now:\[15\div(-4^2)\] See. It is simpler now. \[-4^2 = -4\times-4=-16\] Now our equation looks like this: \[15\div-16\] Now for the last step: The answer is -0.9375 But let me check first...I think I am right, but just to be sure, instead of putting the decimal put 15 over 16 which is definitely the right answer.

OpenStudy (kewlgeek555):

I think @PFEH.1999 is wrong because he did the exponent first. I know PEMDAS, but if there is more than one of the same operations, it is left to right. He worked on the paranthesis right to left, which is wrong.

OpenStudy (kewlgeek555):

I doubt it is 12. I know it is \[\frac{ 15 }{ 16 }\]

OpenStudy (anonymous):

yes,you're right,but i just was asking him some questions and i didn't do that,besides that i didn't find any answer :)

OpenStudy (kewlgeek555):

Oh okay. I thought you said the answer to the whole equation was 12. Sorry for the misunderstandment.